# Topological restrictions on relatively Anosov representations

### v1

@@ 5.1. Restrictions on topological dimension of the boundary
In general, if a group $\Gamma$ admits a $k$-Anosov representation into $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$, where $1\leq k<\frac{d}{2}$, then the topological dimension of the boundary $\partial_{\infty}\Gamma$ is at most $d-k-1$ [CT20, Thm. 1.3]. The same argument in fact works for relatively $k$-Anosov representations.
@@
This follows from [KB02, Thm. 4.4]. The result is stated there for Gromov boundaries of hyperbolic groups, but also applies to the Bowditch boundary of a non-elementary relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ with the same proof, since pairs $(\gamma^{+},\gamma^{-})$ of fixed points of loxodromic elements in any relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ are dense in $(\partial(\Gamma,\operatorname{\mathcal{P}}))^{2}$. ∎
@@
We work as in the proof of [CT20]. Let us fix $x_{0}\in\partial(\Gamma,\mathcal{P})$, a vector $v\in\mathbb{R}^{d}\smallsetminus\xi_{\rho}^{d-k}(x_{0})$ and consider the map
@@
In particular, if $\textup{dim}(\partial(\Gamma,\mathcal{P}))=d-k$, then, arguing as in the proof of [CT20, Thm. 1.3], we conclude that the image of the map $f_{v}$ has to contain an open subset of the affine chart $\mathbb{P}\big{(}\xi_{\rho}^{d-k}(x_{0})\oplus\mathbb{R}v\big{)}\smallsetminus\mathbb{P}(\xi_{\rho}^{d-k}(x_{0}))$ and thus, by Proposition 5.2, we have $\partial(\Gamma,\mathcal{P})\cong S^{d-k}$. In particular, $f_{v}$ is onto and therefore $\mathcal{E}_{\rho,k}$ contains $\mathbb{P}(\mathbb{R}^{d})\smallsetminus\mathbb{P}(\xi_{\rho}^{d-k}(x_{0}))$. Thus, we have $\mathcal{E}_{\rho,k}=\mathbb{P}(\mathbb{R}^{d})$, so we obtain a fibration $\mathbb{P}(\mathbb{R}^{k})\xhookrightarrow{}\mathbb{P}(\mathbb{R}^{d})\twoheadrightarrow S^{d-k}$. By the classification of such fibrations by Adams [Ada60], we deduce that $(d,k)\in\{(2,1),(4,2),(8,4),(16,8)\}$.∎
@@ 5.2. Excluding spheres in the boundary
The upper bound for the cohomological dimension of the domain group of an Anosov representation provided by [CT20, Thm. 1.3] is known to be achieved only in a few cases. The exceptional cases where the bound of Proposition 5.1 is achieved are when either $k\in\{1,2,4\}$, see also the discussion in [CT20, §10]. The case where $k=1$ is achieved when $\Gamma$ is a lattice in $\mathsf{SO}(d,1)$. The case where $k=2$ is achieved is when $\Gamma$ is a lattice in $\mathsf{SU}(d,1)$ and we restrict the natural embedding $\mathsf{SL}_{d+1}(\mathbb{C})\xhookrightarrow{}\mathsf{SL}_{2d+2}(\mathbb{R})$ on $\Gamma<\mathsf{SU}(d,1)$. Finally, the case where $k=4$ is achieved when $\Gamma$ is a lattice in $\operatorname{\mathsf{Sp}}(d,1)$ and we restrict the natural embedding $\mathsf{SL}_{d+1}(\mathbb{H})\xhookrightarrow{}\mathsf{SL}_{4d+4}(\mathbb{R})$ on $\Gamma<\operatorname{\mathsf{Sp}}(d,1)$.
@@
The following theorem, besides being useful later in the paper, offers some further evidence for a positive answer to this question.
@@
Let $d,k\in\operatorname{\mathbb{N}}$ with $d\geq 2k+1$ and $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $(\Gamma,\operatorname{\mathcal{P}})$ be a relatively hyperbolic group whose Bowditch boundary $\partial(\Gamma,\operatorname{\mathcal{P}})$ contains properly a sphere of dimension $r_{\operatorname{\mathbb{K}}}(d-k)-1$ (where $r_{\operatorname{\mathbb{R}}}=1$ and $r_{\operatorname{\mathbb{C}}}=2$). Then there is no relatively $k$-Anosov representation $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{K}})$.
@@
We will prove the statement when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$. The case where $\mathbb{K}=\mathbb{C}$ then follows follows by the real case, since the composition of a $k$-Anosov representation of $(\Gamma,\mathcal{P})$ into $\mathsf{SL}_{d}(\mathbb{C})$, with the standard embedding of $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{C}})\xhookrightarrow{}\operatorname{\mathsf{SL}}_{2d}(\operatorname{\mathbb{R}})$, is a relatively $2k$-Anosov representation into $\operatorname{\mathsf{SL}}_{2d}(\operatorname{\mathbb{R}})$.
Let us set $m=m(d,k):=d-k-1$. By assumption, there is an embedding $f\colon S^{m}\hookrightarrow\partial(\Gamma,\operatorname{\mathcal{P}})$ and $x_{0}\in\partial(\Gamma,\operatorname{\mathcal{P}})\smallsetminus f(S^{m})$. We will argue by contradiction. Suppose that there is a relatively $k$-Anosov representation $\rho\colon(\Gamma,\mathcal{P})\to\mathsf{SL}_{d}(\operatorname{\mathbb{R}})$ with transverse limit maps $(\xi_{\rho}^{k},\xi_{\rho}^{d-k})\colon\partial(\Gamma,\mathcal{P})\rightarrow\mathsf{Gr}_{k}(\mathbb{R}^{d})\times\mathsf{Gr}_{d-k}(\mathbb{R}^{d})$. Note, by Definition 2.8, that the dual representation $\rho^{\ast}\colon(\Gamma,\mathcal{P})\rightarrow\mathsf{SL}_{d}(\mathbb{R})$ is also relatively $k$-Anosov, and
$\xi_{\rho}^{k}(x)=\big{(}\xi_{\rho^{\ast}}^{d-k}(x)\big{)}^{\perp}\quad\forall\ x\in\partial(\Gamma,\mathcal{P}).$ (1)
Let $W_{0}:=\xi_{\rho}^{d-k}(x_{0})\cap\xi_{\rho^{\ast}}^{d-k}(x_{0})$. Since $\xi_{\rho^{\ast}}^{k}(x_{0})\subset\xi_{\rho^{\ast}}^{d-k}(x_{0})$, by (1), we have $\dim(W_{0})=d-2k$. Let us also fix non-zero vectors $v_{1}\in W_{0}$ and $\omega_{1}\in\xi_{\rho}^{k}(x_{0})$. Consider the maps
$\mathsf{F}$ $\colon S^{m}\to\mathbb{P}\big{(}\xi_{\rho^{\ast}}^{d-k}(x_{0})\oplus\operatorname{\mathbb{R}}\omega_{1}\big{)}\smallsetminus\mathbb{P}\big{(}\xi_{\rho^{\ast}}^{d-k}(x_{0})\big{)},$
$\mathsf{G}$ $\colon S^{m}\to\mathbb{P}\big{(}\xi_{\rho}^{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\big{)}\smallsetminus\mathbb{P}(\xi_{\rho}^{k}(x_{0})),$
@@
$\mathsf{F}(x)$ $:=\big{[}\xi_{\rho^{*}}^{k}(f(x))\cap\big{(}\xi_{\rho^{*}}^{d-k}(x_{0})\oplus\operatorname{\mathbb{R}}\omega_{1}\big{)}\big{]},$
$\mathsf{G}(x)$ $:=\big{[}\xi_{\rho}^{d-k}(f(x))\cap\big{(}\xi_{\rho}^{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\big{)}\big{]},$
$\mathsf{H}(x)$ $:=\big{[}\xi_{\rho}^{d-k}(f(x))\cap\big{(}\xi_{\rho}^{k}(f(-x))\oplus\mathsf{G}(-x)\big{)}\big{]}.$
Note that $\mathsf{F}\colon S^{m}\rightarrow\mathbb{P}(\xi_{\rho^{\ast}}^{d-k}(x_{0})\oplus\mathbb{R}\omega_{1})$ and $\mathsf{G}\colon S^{m}\to\mathbb{P}(\xi_{\rho}^{k}(x_{0})\oplus\mathbb{R}v_{1})$ are well-defined and continuous by transversality. In addition, $\mathsf{H}:S^{m}\rightarrow\mathbb{P}(\mathbb{R}^{d})$ is well-defined and continuous since $\xi_{\rho}^{d-k}(f(x))\oplus\xi_{\rho}^{k}(f(-x))=\operatorname{\mathbb{R}}^{d}$ for every $x\in S^{m}$ and $\mathsf{G}(-x)\in\xi_{\rho}^{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\subset\xi_{\rho}^{d-k}(x_{0})$ is transverse to $\xi_{\rho}^{k}(f(-x))$ since $x_{0}\in\partial(\Gamma,\mathcal{P})\smallsetminus f(S^{m})$. Observe also that there are continuous functions $\mathsf{F}_{0}\colon S^{m}\to\xi_{\rho^{\ast}}^{d-k}(x_{0})$ and $\mathsf{G}_{0}\colon S^{m}\to\xi_{\rho}^{k}(x_{0})$ such that
@@
Now we are going to construct a lift for the map $\mathsf{H}$. First, note that for every $x\in S^{m}$ a lift of $\mathsf{H}(x)\in\mathbb{P}(\xi_{\rho}^{d-k}(f(x)))$ can be chosen to be of the form $\mathsf{R}(x)+\widetilde{\mathsf{G}}(-x)$ for some $\mathsf{R}(x)\in\xi_{\rho}^{k}(f(-x))$, since $\xi_{\rho}^{k}(f(-x))\cap\xi_{\rho}^{d-k}(f(x))=\{\bf 0\}$. Thus, there are maps $\widetilde{\mathsf{H}}\colon S^{m}\to\operatorname{\mathbb{R}}^{d}\smallsetminus\{\bf 0\}$ and $\mathsf{R}\colon S^{m}\to\operatorname{\mathbb{R}}^{d}$ with the property that
@@
is a lift of $\mathsf{H}$, $\widetilde{\mathsf{H}}(x)\in\xi_{\rho}^{d-k}(f(x))$ and $\mathsf{R}(x)\in\xi_{\rho}^{k}(f(-x))$ for every $x\in S^{m}$.
Claim: The maps $\widetilde{\mathsf{H}}$ and $\mathsf{R}$ are well-defined and continuous. We first observe that $\mathsf{R}(x)\in\xi_{\rho}^{k}(f(-x))$ is the unique vector such that $\mathsf{R}(x)+\mathsf{G}_{0}(-x)+v_{1}$ is a lift of $\mathsf{H}(x)\in\mathbb{P}(\xi_{\rho}^{d-k}(f(x)))$. Indeed, if $\mathsf{R}^{\prime}(x)\in\xi_{\rho}^{k}(f(-x))$ is another vector such that $[\mathsf{R}^{\prime}(x)+G_{0}(-x)+v_{1}]=\mathsf{H}(x)$, then
$\mathsf{R}(x)-\mathsf{R}^{\prime}(x)\in\xi_{\rho}^{k}(f(-x))\cap\xi_{\rho}^{d-k}(f(x))=\{\bf 0\}$
@@
Now in order to prove that $\mathsf{R}\colon S^{m}\rightarrow\mathbb{R}^{d}$ is continuous we first observe that the map $x\mapsto\left\|\mathsf{R}(x)\right\|$ is bounded. If not, there is a sequence $(y_{n})_{n\in\mathbb{N}}\subset S^{m}$ with $\lim_{n}y_{n}=y_{0}$ and $\lim_{n}||\mathsf{R}(y_{n})||=\infty$. Since $x\mapsto\left\|\widetilde{\mathsf{G}}_{0}(x)+v_{1}\right\|$ is continuous and bounded, up to a passing to a subsequence, we have
@@
where both limits exist and are non-zero vectors of $S^{d-1}$. Moreover, since $\frac{1}{||\mathsf{R}(y_{n})||}\mathsf{R}(y_{n})\in\xi_{\rho}^{k}(f(-y_{n}))$ for every $n\in\mathbb{N}$, we have $\lim_{n}\frac{1}{||\mathsf{R}(y_{n})||}\mathsf{R}(y_{n})\in S^{1}(\xi_{\rho}^{k}(f(-y_{0})))$. Similarly, $\frac{1}{||\mathsf{R}(y_{n})||}\widetilde{\mathsf{H}}(y_{n})\in\xi_{\rho}^{d-k}(f(y_{n}))$ for every $n$, so $\lim_{n}\frac{1}{||\mathsf{R}(y_{n})||}\widetilde{\mathsf{H}}(y_{n})\in S^{1}(\xi_{\rho}^{d-k}(f(y_{0})))$. However, by (2), this contradicts that by transversality we have $\xi_{\rho}^{k}(f(-y_{0}))\cap\xi_{\rho}^{d-k}(f(y_{0}))=\{\bf 0\}$. It follows that the map $x\mapsto\mathsf{R}(x)$ is bounded.
Now it is a formality to check that $\mathsf{R}$ and $\widetilde{\mathsf{H}}$ are continuous. Let $(z_{n})_{n\in\mathbb{N}}\subset S^{m}$ be an arbitrary sequence with $\lim_{n}z_{n}=z_{0}$ and $\mathsf{R}_{\infty}\in\xi_{\rho}^{k}(f(-z_{0}))$ an accumulation point of $(\mathsf{R}(z_{n}))_{n\in\mathbb{N}}$. Then, note that $\mathsf{R}_{\infty}+\mathsf{G}_{0}(-z_{0})+v_{1}\neq 0$ is a lift of $\mathsf{H}(z_{0})\in\mathbb{P}(\xi_{\rho}^{d-k}(f(z_{0})))$. Hence there is $\lambda\in\operatorname{\mathbb{R}}^{\ast}$ such that
@@
Now note that $\mathsf{R}_{\infty}\in\xi_{\rho}^{k}(f(-z_{0}))$, $\mathsf{R}(z_{0})\in\xi_{\rho}^{k}(f(-z_{0}))$, $\mathsf{G}_{0}(-z_{0})+v_{1}\in\xi_{\rho}^{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}$ and
$\xi_{\rho}^{k}(f(-z_{0}))\cap(\xi_{\rho}^{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1})\subset\xi_{\rho}^{k}(f(-z_{0}))\cap\xi_{\rho}^{d-k}(x_{0})=\{\bf 0\},$
so we have $\lambda=1$ and $\mathsf{R}_{\infty}=\mathsf{R}(z_{0})$. Finally, the claim follows and $\widetilde{\mathsf{H}}$ and $\mathsf{R}$ are well-defined and continuous.
@@
$\operatorname{\mathbb{R}}^{d}=\xi_{\rho}^{k}(x_{0})\oplus\xi_{\rho^{\ast}}^{d-k}(x_{0})$ (3)

### v2

@@ 5.1. Restrictions on topological dimension of the boundary
In general, if a group $\Gamma$ admits a $k$-Anosov representation into $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$, where $1\leq k<\frac{d}{2}$, then the topological dimension of the boundary $\partial_{\infty}\Gamma$ is at most $d-k-1$ [CanaryTsouvalas, Thm. 1.3]. The same argument in fact works for relatively $k$-Anosov representations.
@@
This follows from [KapoBen, Thm. 4.4]. The result is stated there for Gromov boundaries of hyperbolic groups, but also applies to the Bowditch boundary of a non-elementary relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ with the same proof, since pairs $(\gamma^{+},\gamma^{-})$ of fixed points of loxodromic elements in any relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ are dense in $(\partial(\Gamma,\operatorname{\mathcal{P}}))^{2}$. ∎
@@
We work as in the proof of [CanaryTsouvalas]. Let us fix $x_{0}\in\partial(\Gamma,\mathcal{P})$, a vector $v\in\mathbb{R}^{d}\smallsetminus\xi_{\rho}^{d-k}(x_{0})$ and consider the map
@@
In particular, if $\textup{dim}(\partial(\Gamma,\mathcal{P}))=d-k$, then, arguing as in the proof of [CanaryTsouvalas, Thm. 1.3], we conclude that the image of the map $f_{v}$ has to contain an open subset of the affine chart $\mathbb{P}\big{(}\xi_{\rho}^{d-k}(x_{0})\oplus\mathbb{R}v\big{)}\smallsetminus\mathbb{P}(\xi_{\rho}^{d-k}(x_{0}))$ and thus, by Proposition 5.2, we have $\partial(\Gamma,\mathcal{P})\cong S^{d-k}$. In particular, $f_{v}$ is onto and therefore $\mathcal{E}_{\rho,k}$ contains $\mathbb{P}(\mathbb{R}^{d})\smallsetminus\mathbb{P}(\xi_{\rho}^{d-k}(x_{0}))$. Thus, we have $\mathcal{E}_{\rho,k}=\mathbb{P}(\mathbb{R}^{d})$, so we obtain a fibration $\mathbb{P}(\mathbb{R}^{k})\xhookrightarrow{}\mathbb{P}(\mathbb{R}^{d})\twoheadrightarrow S^{d-k}$. By the classification of such fibrations by Adams [Adams], we deduce that $(d,k)\in\{(2,1),(4,2),(8,4),(16,8)\}$.∎
@@ 5.2. Excluding spheres in the boundary
The upper bound for the cohomological dimension of the domain group of an Anosov representation provided by [CanaryTsouvalas, Thm. 1.3] is known to be achieved only in a few cases. The exceptional cases where the bound of Proposition 5.1 is achieved are when either $k\in\{1,2,4\}$, see also the discussion in [CanaryTsouvalas, §10]. The case where $k=1$ is achieved when $\Gamma$ is a lattice in $\mathsf{SO}(d,1)$. The case where $k=2$ is achieved is when $\Gamma$ is a lattice in $\mathsf{SU}(d,1)$ and we restrict the natural embedding $\mathsf{SL}_{d+1}(\mathbb{C})\xhookrightarrow{}\mathsf{SL}_{2d+2}(\mathbb{R})$ on $\Gamma<\mathsf{SU}(d,1)$. Finally, the case where $k=4$ is achieved when $\Gamma$ is a lattice in $\operatorname{\mathsf{Sp}}(d,1)$ and we restrict the natural embedding $\mathsf{SL}_{d+1}(\mathbb{H})\xhookrightarrow{}\mathsf{SL}_{4d+4}(\mathbb{R})$ on $\Gamma<\operatorname{\mathsf{Sp}}(d,1)$.
@@
Theorem 1.6 offers some further evidence for a positive answer to this question. It will be a consequence of the following more general theorem, since the limit maps associated to a relatively Anosov representation yield a transverse subset. We remark that the more general theorem also applies to e.g. $\operatorname{\mathsf{P}}_{k}$-transverse subgroups of $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{K}})$, as defined by Canary–Zhang–Zimmer [CZZ].
@@
Let $d,k\in\operatorname{\mathbb{N}}$ with $d\geq 2k+1$ and $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. There is no transverse set
$\Lambda\subset\operatorname{\mathsf{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\times\operatorname{\mathsf{Gr}}_{d-k}(\operatorname{\mathbb{K}}^{d})$
which properly contains a sphere of dimension $r_{\operatorname{\mathbb{K}}}(d-k)-1$ (where $r_{\operatorname{\mathbb{R}}}=1$ and $r_{\operatorname{\mathbb{C}}}=2$).
@@
There will be several cases, depending on $d$ and $k$; in all cases, we will argue by contradiction. The following case is more straightforward, and we will reuse the main idea in its proof in later results.
Lemma 5.4.
Suppose $d=2k+1$ and $k\in\mathbb{N}$ is odd. Suppose $\Lambda\subset\operatorname{\mathsf{Gr}}_{k}(\operatorname{\mathbb{R}}^{d})\times\operatorname{\mathsf{Gr}}_{d-k}(\operatorname{\mathbb{R}}^{d})$ is a transverse set. Then $\Lambda$ cannot properly contain a $k$-sphere.
Proof.
We argue by contradiction. Suppose that there is a transverse set $\Lambda\subset\operatorname{\mathsf{Gr}}_{k}(\operatorname{\mathbb{R}}^{d})\times\operatorname{\mathsf{Gr}}_{d-k}(\operatorname{\mathbb{R}}^{d})$, an embedding $f\colon S^{k}\hookrightarrow\Lambda$ and $x_{0}\in\Lambda\smallsetminus f(S^{k})$.
Let $\pi_{k}\colon\Lambda\to\operatorname{\mathsf{Gr}}_{k}(\operatorname{\mathbb{R}}^{d})$ and $\pi_{k+1}\colon\Lambda\to\operatorname{\mathsf{Gr}}_{k+1}(\operatorname{\mathbb{R}}^{d})$ be the natural projections. By transversality, these are homeomorphisms.
Note that since $x_{0}\in\partial(\Gamma,P)\smallsetminus f(S^{k})$, by transversality, the map $\tau_{k}^{+}\circ\pi_{k}\circ f\colon S^{k}\to\mathbb{P}(\wedge^{k}\operatorname{\mathbb{R}}^{d})$ avoids the hyperplane $\tau_{k}^{-}(\pi_{d-k}(x_{0}))\subset\wedge^{k}\mathbb{R}^{2k+1}$. Thus, we obtain a continuous lift $\widetilde{\xi}_{k}\colon S^{k}\to\mathbb{S}(\wedge^{k}\operatorname{\mathbb{R}}^{d})$ of $\tau_{k}^{+}\circ\pi_{k}\circ f$, where $\mathbb{S}(\cdot)$ denotes the unit sphere in the vector space. In addition, the continuous map $\mathsf{F}_{1}\colon S^{k}\to\mathbb{P}(\pi_{k+1}(x_{0}))\smallsetminus\mathbb{P}(\pi_{k}(x_{0}))$ defined by
$\mathsf{F}_{1}(x)=\big{[}\pi_{k+1}(f(x))\cap\pi_{k+1}(x_{0})\big{]}$
admits a continuous lift, i.e. there is $v_{0}\in\pi_{k+1}(x_{0})\smallsetminus\pi_{k}(x_{0})$ and a continuous map $\mathsf{F}_{1}^{\prime}\colon S^{k}\to\pi_{k}(x_{0})\cong\operatorname{\mathbb{R}}^{k}$ such that $[\mathsf{F}_{1}^{\prime}(x)+v_{0}]=[\mathsf{F}_{1}(x)]$ for every $x\in S^{k}$.
By transversality for $x,y\in S^{k}$ distinct we have that $\pi_{k+1}(f(x))=\pi_{k}(f(x))\oplus\mathbb{R}\mathsf{F}_{1}(x)$ and
$\operatorname{\mathbb{R}}^{d}=\pi_{k}(f(x))\oplus\pi_{k}(f(y))\oplus\mathbb{R}\mathsf{F}_{1}(x),$
so we obtain a continuous, non-vanishing map $\mathcal{H}\colon(S^{k})^{(2)}\to\operatorname{\mathbb{R}}^{\ast}$ given by
Now $\mathsf{F}_{1}^{\prime}$ is a continuous map from the $k$-sphere to an affine chart in projective $k$-space, and hence cannot be injective. Hence there are $x_{1},y_{1}\in S^{k}$ distinct with $\mathsf{F}_{1}^{\prime}(x_{1})=\mathsf{F}_{1}^{\prime}(y_{1})$. But then, since $k\in\mathbb{N}$ is odd, we have
$=-\mathcal{H}(x_{1},y_{1}).$
This contradicts the fact that $\operatorname{\mathcal{H}}$ is a non-zero continuous map. ∎
@@
We will prove the statement when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$. The case where $\mathbb{K}=\mathbb{C}$ then follows by the real case, since this says there is no transverse subset of $\operatorname{\mathsf{Gr}}_{2k}(\operatorname{\mathbb{R}}^{2d})\times\operatorname{\mathsf{Gr}}_{2d-2k}(\operatorname{\mathbb{R}}^{2d})$ properly containing a $(2d-2k-1)$-sphere, and $\operatorname{\mathsf{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})\subset\operatorname{\mathsf{Gr}}_{2k}(\operatorname{\mathbb{R}}^{2d})$.
We argue by contradiction.
Set $m=m(d,k):=d-k-1$. Suppose there is a transverse subset $\Lambda\subset\mathsf{Gr}_{k}(\mathbb{R}^{d})\times\mathsf{Gr}_{d-k}(\mathbb{R}^{d})$ and an embedding $f\colon S^{m}\hookrightarrow\Lambda$ and $x_{0}\in\Lambda\smallsetminus f(S^{m})$. If $d=2k+1$ and $k$ is odd, then we are done by Lemma 5.4.
The proof of the other cases will be more involved.
Let $W_{0}:=\pi_{d-k}(x_{0})\cap(\pi_{k}(x_{0}))^{\perp}$. We have $\dim(W_{0})=d-2k$. Let us also fix non-zero vectors $v_{1}\in W_{0}$ and $\omega_{1}\in\pi_{k}(x_{0})$. Consider the maps
$\mathsf{F}$ $\colon S^{m}\to\mathbb{P}\big{(}(\pi_{k}(x_{0}))^{\perp}\oplus\operatorname{\mathbb{R}}\omega_{1}\big{)}\smallsetminus\mathbb{P}\big{(}(\pi_{k}(x_{0}))^{\perp}\big{)},$
$\mathsf{G}$ $\colon S^{m}\to\mathbb{P}\big{(}\pi_{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\big{)}\smallsetminus\mathbb{P}(\pi_{k}(x_{0})),$
@@
$\mathsf{F}(x)$ $:=(\pi_{d-k}(f(x)))^{\perp}\cap\big{(}(\pi_{k}(x_{0}))^{\perp}\oplus\operatorname{\mathbb{R}}\omega_{1}\big{)}\big{]},$
$\mathsf{G}(x)$ $:=\big{[}\pi_{d-k}(f(x))\cap\big{(}\pi_{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\big{)}\big{]},$
$\mathsf{H}(x)$ $:=\big{[}\pi_{d-k}(f(x))\cap\big{(}\pi_{k}(f(-x))\oplus\mathsf{G}(-x)\big{)}\big{]}.$
Note that $\mathsf{F}\colon S^{m}\rightarrow\mathbb{P}(\xi_{\rho^{\ast}}^{d-k}(x_{0})\oplus\mathbb{R}\omega_{1})$ and $\mathsf{G}\colon S^{m}\to\mathbb{P}(\pi_{k}(x_{0})\oplus\mathbb{R}v_{1})$ are well-defined and continuous by transversality. In addition, $\mathsf{H}:S^{m}\rightarrow\mathbb{P}(\mathbb{R}^{d})$ is well-defined and continuous since $\pi_{d-k}(f(x))\oplus\pi_{k}(f(-x))=\operatorname{\mathbb{R}}^{d}$ for every $x\in S^{m}$ and $\mathsf{G}(-x)\in\pi_{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}\subset\pi_{d-k}(x_{0})$ is transverse to $\pi_{k}(f(-x))$ since $x_{0}\in\partial(\Gamma,\mathcal{P})\smallsetminus f(S^{m})$. Observe also that there are continuous functions $\mathsf{F}_{0}\colon S^{m}\to(\pi_{k}(x_{0}))^{\perp}$ and $\mathsf{G}_{0}\colon S^{m}\to\pi_{k}(x_{0})$ such that
@@
Now we are going to construct a lift for the map $\mathsf{H}$. First, note that for every $x\in S^{m}$ a lift of $\mathsf{H}(x)\in\mathbb{P}(\pi_{d-k}(f(x)))$ can be chosen to be of the form $\mathsf{R}(x)+\widetilde{\mathsf{G}}(-x)$ for some $\mathsf{R}(x)\in\pi_{k}(f(-x))$, since $\pi_{k}(f(-x))\cap\pi_{d-k}(f(x))=\{\bf 0\}$. Thus, there are maps $\widetilde{\mathsf{H}}\colon S^{m}\to\operatorname{\mathbb{R}}^{d}\smallsetminus\{\bf 0\}$ and $\mathsf{R}\colon S^{m}\to\operatorname{\mathbb{R}}^{d}$ with the property that
@@
is a lift of $\mathsf{H}$, $\widetilde{\mathsf{H}}(x)\in\pi_{d-k}(f(x))$ and $\mathsf{R}(x)\in\pi_{k}(f(-x))$ for every $x\in S^{m}$.
Claim: The maps $\widetilde{\mathsf{H}}$ and $\mathsf{R}$ are well-defined and continuous. If $\mathsf{R}^{\prime}(x),\mathsf{R}^{\prime}(x)\in\pi_{k}(f(-x))$ are two vectors such that $[\mathsf{R}(x)+\mathsf{G}_{0}(-x)+v_{1}]=[\mathsf{R}^{\prime}(x)+\mathsf{G}_{0}(-x)+v_{1}]=\mathsf{H}(x)$, then
$\mathsf{R}(x)-\mathsf{R}^{\prime}(x)\in\pi_{k}(f(-x))\cap\pi_{d-k}(f(x))=\{\bf 0\}$
@@
Now in order to prove that $\mathsf{R}\colon S^{m}\rightarrow\mathbb{R}^{d}$ is continuous we first observe that the map $x\mapsto\left\|\mathsf{R}(x)\right\|$ is bounded. If not, there is a sequence $(y_{n})_{n\in\mathbb{N}}\subset S^{m}$ with $\lim_{n}y_{n}=y_{0}$ and $\lim_{n}||\mathsf{R}(y_{n})||=\infty$. Since $x\mapsto\left\|{\mathsf{G}}_{0}(x)+v_{1}\right\|$ is continuous and bounded, up to a passing to a subsequence, we have
@@
where both limits exist and are non-zero vectors of $S^{d-1}$. Moreover, since $\frac{1}{||\mathsf{R}(y_{n})||}\mathsf{R}(y_{n})\in\pi_{k}(f(-y_{n}))$ for every $n\in\mathbb{N}$, we have $\lim_{n}\frac{1}{||\mathsf{R}(y_{n})||}\mathsf{R}(y_{n})\in\mathbb{S}(\pi_{k}(f(-y_{0})))$. Similarly, $\frac{1}{||\mathsf{R}(y_{n})||}\widetilde{\mathsf{H}}(y_{n})\in\pi_{d-k}(f(y_{n}))$ for every $n$, so $\lim_{n}\frac{1}{||\mathsf{R}(y_{n})||}\widetilde{\mathsf{H}}(y_{n})\in\mathbb{S}(\pi_{d-k}(f(y_{0})))$. However, by (1), this contradicts that by transversality we have $\pi_{k}(f(-y_{0}))\cap\pi_{d-k}(f(y_{0}))=\{\bf 0\}$. It follows that the map $x\mapsto\mathsf{R}(x)$ is bounded.
Now it is a formality to check that $\mathsf{R}$ and $\widetilde{\mathsf{H}}$ are continuous. Let $(z_{n})_{n\in\mathbb{N}}\subset S^{m}$ be an arbitrary sequence with $\lim_{n}z_{n}=z_{0}$ and $\mathsf{R}_{\infty}\in\pi_{k}(f(-z_{0}))$ an accumulation point of $(\mathsf{R}(z_{n}))_{n\in\mathbb{N}}$. Then, note that $\mathsf{R}_{\infty}+\mathsf{G}_{0}(-z_{0})+v_{1}\neq 0$ is a lift of $\mathsf{H}(z_{0})\in\mathbb{P}(\pi_{d-k}(f(z_{0})))$. Hence there is $\lambda\in\operatorname{\mathbb{R}}^{\ast}$ such that
@@
Now note that $\mathsf{R}_{\infty}\in\pi_{k}(f(-z_{0}))$, $\mathsf{R}(z_{0})\in\pi_{k}(f(-z_{0}))$, $\mathsf{G}_{0}(-z_{0})+v_{1}\in\pi_{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1}$ and
$\pi_{k}(f(-z_{0}))\cap(\pi_{k}(x_{0})\oplus\operatorname{\mathbb{R}}v_{1})\subset\pi_{k}(f(-z_{0}))\cap\pi_{d-k}(x_{0})=\{\bf 0\},$
so we have $\lambda=1$ and $\mathsf{R}_{\infty}=\mathsf{R}(z_{0})$. Finally, the claim follows and $\widetilde{\mathsf{H}}$ and $\mathsf{R}$ are well-defined and continuous. ∎
@@
$\operatorname{\mathbb{R}}^{d}=\pi_{k}(x_{0})\oplus(\pi_{k}(x_{0}))^{\perp}$ (2)